17
Feb
2011
03:55

The earlier issue involving colo customers has been traced to a packet storm, originating from a specific customer machine which has been isolated. The effect on other customers was due to a lack of resources processing a very high number of new sessions – the FireBrick team will be investigating ways to provide better isolation between customers. In the meantime, the network has now been running stably since 03:30.

Posted by Kevin Hones

17
Feb
2011
02:06

We are currently experiencing problems with our colo customer edge router in Bluesquare 2. Apologies for the break in service – investigations are underway.

Posted by admin

14
Feb
2011
10:59

Latest information from Fluidata “This issue is now under control, having isolated a switch from the network on this VLAN. We are still investigating why this switch was behaving incorrectly and will provide more detailed information later on.”

We apologise for the disruption to some services.

Posted by admin

14
Feb
2011
10:00

Fluidata are reporting they “lost a fibre span between Telehouse North and London Hosting Centre at 08:33 this morning. This has been logged to our dark fibre provider for investigation. There is some instability on the wholesale platform as a result, and our engineers are currently investigating this.” We will provide updates as they become available.

Posted by admin

04
Feb
2011
00:28

We suffered a power outage in our “MHB” datacentre at 23:09 yesterday evening. This resulted in connectivity issues for some Be ADSL lines, most bonded ADSL customers and legacy equipment in London. The datacentre are investigating the cause, but power has now been restored. Apologies for this outage.

Posted by Kevin Hones

Current Status

    GoodNo current issues

Status Archive